# Euler systems and relative Satake isomorphism

### v1

@@ 3.5. Abstract Euler system I
To state the abstract Euler system relation, we need to specify the Hecke operator. Our approach depends very little on this choice.
@@
[…]
Example 3.12.
Let $V$ be a minuscule representation of $\check{G}$, then the function $t\mapsto\det(1-q^{-\frac{1}{2}}t|V)$ belongs to $\mathbb{C}[A^{*}]^{W}$. The Hecke operator associated to it by the Satake transform lies in $\mathbb{Z}_{(q)}[A^{*}]$, so it is $\mathbf{X}$-integral for all $\mathbf{X}$. More generally, replacing $t$ by $tX$ gives a polynomial $\mathcal{H}_{V}(X)$, all of whose coefficients are $\mathbf{X}$-integral. This is the Hecke polynomial considered in [BR94, §6].
@@
Let $\mathcal{L}$ be an $\widetilde{\mathbf{X}}$-integral Hecke operator, then there exists $\Phi_{1}\in C_{c}^{\infty}({\widetilde{X}},\mathbb{Z}[q^{-1}])^{K^{1}}$ such that
$\operatorname{Tr}_{K^{0}}^{K^{1}}\Phi_{1}=\mathcal{L}\cdot\mathbf{1}[\mathbf{X}(\mathcal{O})],$
@@
By Proposition 3.4, if this coefficient is not 1, then $x$ lies on a wall of type T. Proposition 3.9 applied to $\widetilde{\mathbf{X}}$ shows that this divisibility requirement is automatic. ∎
4. Global picture
@@ 4.1. Motivic theta series
$\delta:=\prod_{v\notin S}\mathbf{1}[\mathbf{X}(\mathcal{O}_{F_{v}})]\in C_{c}^{\infty}(\mathbf{X}(\mathbb{A}^{S}),\mathbb{Z}_{p})$
@@
The basic element $z\in\mathcal{M}$ is defined to be $\Theta(\delta)$.
@@ 4.2. Abstract Euler system II
Let $\mathscr{L}$ be the set of places of $F$ away from $S$ where the hypotheses of Section 2 are satisfied for $\widetilde{\mathbf{X}}$. This in particular requires $\mathbf{T}$ to be split at those places. For each $\ell\in\mathscr{L}$, let $\mathcal{H}_{\ell}$ be an $\widetilde{\mathbf{X}}$-integral Hecke operator in the group $\widetilde{\mathbf{G}}(F_{\ell})$ (Definition 3.11). They give rise to the following abstract Euler system.
@@
For each $\ell\in\mathscr{L}$, let $\delta^{(\ell)}$ be the element $\Phi_{1}$ from Proposition 3.13. Define $\delta[\mathfrak{m}]$ to be $\delta$, except the local component at each $\ell|\mathfrak{m}$ are replaced by $\delta^{(\ell)}$. The theorem follows from the $\widetilde{\mathbf{G}}(\mathbb{A}^{\infty})$-equivariance of $\Theta$. ∎
@@
$\mathcal{H}(\widetilde{\mathbf{G}}(F_{\ell}),\mathbb{C})\simeq\mathcal{H}(\mathbf{G}(F_{\ell}),\mathbb{C})[\mathtt{T}^{\pm 1}].$
Here, $\mathtt{T}$ is the indicator function of $\varpi\mathcal{O}_{F_{\ell}}^{\times}\subseteq\mathbf{T}(F_{\ell})$. Moreover, a Hecke operator is $\widetilde{\mathbf{X}}$-integral if and only if all of its coefficients on the right hand side are $\mathbf{X}$-integral. In particular, if $V$ is a minuscule representation of $\check{G}$, then we may use the Hecke polynomial $\mathcal{H}_{V}(\mathtt{T})$ defined in Example 3.12.
@@
Let the setting be as in Theorem 4.5, with $\mathcal{M}$ given by equation (9). Let $V$ be a minuscule representation of $\check{G}$. Then there exists a collection of elements
@@
whenever $\mathfrak{m},\mathfrak{m}\ell\in\mathscr{R}$.
@@
    Projection to Galois representations: Kottwitz’s conjecture gives a concrete description of the cohomology group $\mathrm{H}^{d-1}(\mathrm{Sh}_{\mathbf{G}/\bar{E}},\mathbb{L}\otimes_{\mathbb{Z}_{p}}\mathbb{Q}_{p})$. In the case considered in Corollary 1.2, it gives a projection from this group to the Galois representation $\rho_{\pi}$. Moreover, the Hecke polynomial specialized to the Satake parameters of $\pi$ agrees with the characteristic polynomial of $\rho_{\pi}$.
@@ 4.4. Special cycles
Under the above assumptions, the discussions of [LS24, §2.4] carries over verbatim, and Proposition 2.7 of op. cit. gives a motivic theta series.
@@
Let $d=\dim\mathrm{Sh}_{\mathbf{H}}$. Let $\mathbb{L}$ be the $\mathbb{Z}_{p}$-local system on $\mathrm{Sh}_{\mathbf{G}}$ attached to $\xi$. The diagonal cycle construction defines an $\mathbf{X}$-theta series
$\Theta:C_{c}^{\infty}(\mathbf{X}(\mathbb{A}^{p\infty}),\mathbb{Z}_{p})\to\mathrm{H}_{\mathrm{cont}}^{2d}(\mathrm{Sh}_{\mathbf{G}},\mathbb{L}(d))$
@@
We now explain the content of Corollary 4.6 in a few cases. In each case, we need to specify the data of an embedding of reductive groups $\iota:\mathbf{H}\hookrightarrow\mathbf{G}$ and a character $\nu:\mathbf{H}\to\mathrm{U}(1)$. The latter is used to specify the $\mathrm{U}(1)$-bundle over $\mathbf{X}$. In all cases, the local spherical variety appears in the table [Sak13, Appendix A], so they have been verified to satisfy the combinatorial conditions alluded to in §3.3.2.
@@ 4.4.1. Gan–Gross–Prasad [LS24]
The intersection of $\mathbf{H}$ with a good Borel $\mathbf{B}$ is trivial, so the bundle is automatically combinatorially trivial. Take $V$ to be the standard tensor product representation of degree $n(n+1)$, then Corollary 4.6 recovers [LS24, Proposition 5.4], bypassing the explicit matrix calculations in §4 of op. cit.
@@ 4.4.2. Friedberg–Jacquet [GS23]
Let $\mathtt{V}_{2n}$ be a Hermitian space with signature $(1,2n-1)$ at one archimedean place and $(0,2n)$ at the other archimedean places. Let $\mathtt{W}$ be a totally definite subspace of dimension $n$, and let $\mathtt{W}^{\perp}$ be its dual. Consider the setting
@@
At a split place, the local picture is $\mathrm{GL}_{n}\times\mathrm{GL}_{n}\backslash\mathrm{GL}_{2n}$. For an appropriate choice of the Borel subgroup $\mathbf{B}$, the intersection $\mathbf{A}\cap\mathbf{H}$ has the form
@@
Take $V$ to be the standard representation of degree $2n$, then Corollary 4.6 recovers the tame part of the split anticyclotomic Euler system for symplectic representations constructed by Graham and Shah [GS23, Proposition 9.27] without assumptions on the prime $p$.
4.4.3. Cornut’s Euler system [Cor18]
Let $F$ be a totally real field. Let $\mathtt{V}$ be a quadratic space over $F$ of signature $(2,2n-1)$ at one archimedean place and istotropic at the other archimedean places. Suppose $\mathtt{V}$ splits over $E$, then there is an $E$-Hermitian $F$-hyperplane $\mathtt{W}\subseteq\mathtt{V}$, unique up to $\mathbf{G}$-translation. Let
$\mathbf{H}=\mathrm{U}(\mathtt{W}),\quad\mathbf{G}=\mathrm{SO}(\mathtt{V})$
The definition gives an inclusion $\iota:\mathbf{H}\hookrightarrow\mathbf{G}$. The character $\nu$ is again the determinant.
At a split place $\ell$, the local picture is $\mathrm{GL}_{n}\backslash\mathrm{SO}_{2n+1}$. The combinatorially trivial condition is automatic. Take $V$ to be the standard representation of degree $2n$. Corollary 4.6 then recovers a split anticyclotomic Euler system for representations of $\mathrm{SO}(\mathtt{V})$, where the “Euler factor” is the one corresponding to the standard $L$-function. This is the split analogue of the one constructed by Cornut [Cor18].
Remark 4.10.
In this case, $\mathbf{A}=\mathbf{A}_{\mathbf{X}}$, but the little Weyl group $W_{X}$ is smaller, so we do not have multiplicity one [Sak08, Theorem 1.2.1]. This observation also shows that $\mathbf{X}$ is not wavefront, which suggests that a representation $\pi$ which is distinguished by $\mathbf{X}$ is necessarily not stable. If this is the case, then Kottwitz’s conjecture predicts that the Galois representation appearing in the orthogonal Shimura variety is not the standard one. We hope further automorphic studies of the setting will shed light on this question.
@@ 4.5. Eisenstein classes
If $\mathbf{H}=\mathrm{GL}_{2}$ and $V$ is its standard two-dimensional representation, then Example 4.3 gives a $V$-theta series as above, with $i=j=1$. This gives rise to the tame parts of the following Euler systems, with the caveat that we need the function-level statements of [SW22] in mixed characteristics.

### v2

@@ 3.5. Abstract Euler system I
We now combine the results from the previous subsections together to produce an abstract tame norm relation.
@@
In the set-up of Propositions 3.11, 3.14, or 3.16, we say that $\mathcal{L}$ is the Hecke operator attached to $\mathbf{X}$. Recall that if $\mathbf{X}$ is strongly tempered, then $\mathcal{L}$ is the denominator of the unramified Plancherel formula for $\mathbf{X}$.
In the rest of this article, we tacitly assume that when this terminology is applied, the relevant pair $(\mathbf{G},\mathbf{X})$ satisfies the conditions in one of the propositions listed. This definition is rather ad hoc beyond the strongly tempered case. One candidate for the general definition is the piece corresponding to $S_{X}$ in the notation of [BZSV24, §9.1], but we will not pursue this further in this paper.
@@
Let $\tilde{\mathbf{X}}\to\mathbf{X}$ be a combinatorially trivial $\mathbb{G}_{m}$-bundle. Let $\mathcal{L}$ be the Hecke operator attached to $\tilde{\mathbf{X}}$, then there exists a function
$\Phi_{1}\in C_{c}^{\infty}({\widetilde{X}},\mathbb{Z}[q^{-1}])^{K^{1}}$
such that
$\operatorname{Tr}_{K^{0}}^{K^{1}}\Phi_{1}=\mathcal{L}\cdot\mathbf{1}[\widetilde{\mathbf{X}}(\mathcal{O})],$
@@
By Proposition 3.4 (applied after a quadratic extension if $\mathbf{G}$ is not split), if this coefficient is not 1, then $x$ lies on a wall of type T. The three propositions cited in the statement applied to $\widetilde{\mathbf{X}}$ shows that this divisibility requirement is satisfied. ∎
Remark 3.19.
  1. (1)
    In all cases, the Hecke operator defined gives rise to the $L$-value attached to the spherical variety $\mathbf{X}$, in the sense of [Sak13, Definition 7.2.3]. This lends credence to the idea the Euler systems, including the tame norm relations, should the “arithmetic shape of zeta functions” [Kat04]. It also suggests further subtleties in non-wavefront cases, such as the one investigated by Cornut [Cor18].
  2. (2)
[…]
4. Construction of Euler systems
@@ 4.1. Motivic theta series
$\Phi_{0}:=\prod_{v\notin S}\mathbf{1}[\mathbf{X}(\mathcal{O}_{F_{v}})]\in C_{c}^{\infty}(\mathbf{X}(\mathbb{A}^{S}),\mathbb{Z}_{p})$
@@
The basic element $z\in\mathcal{M}$ is defined to be $\Theta(\Phi_{0})$.
@@ 4.2. Abstract Euler system II
Let $\mathscr{L}$ be the set of places of $F$ away from $S$ where the hypotheses of Section 2 are satisfied for $\widetilde{\mathbf{X}}$. This in particular requires $\mathbf{T}$ to be split at those places. For each $\ell\in\mathscr{L}$, let $\mathcal{H}_{\ell}$ be the Hecke operator attached to $\mathbf{X}$ (Definition 3.17). They give rise to the following abstract Euler system.
@@
For each $\ell\in\mathscr{L}$, let $\Phi^{(\ell)}$ be the element $\Phi_{1}$ from Proposition 3.18 applied to the completion $F_{\ell}$. Define $\Phi[\mathfrak{m}]$ the same way as $\Phi_{0}$, except the local component at each $\ell|\mathfrak{m}$ are replaced by $\Phi^{(\ell)}$. The theorem follows from the $\widetilde{\mathbf{G}}(\mathbb{A}^{\infty})$-equivariance of $\Theta$. ∎
@@
(11) $\mathcal{H}(\widetilde{\mathbf{G}}(F_{\ell}),\mathbb{C})\simeq\mathcal{H}(\mathbf{G}(F_{\ell}),\mathbb{C})[\mathtt{T}^{\pm 1}].$
Here, $\mathtt{T}$ is the indicator function of $\varpi\mathcal{O}_{F_{\ell}}^{\times}\subseteq\mathbf{T}(F_{\ell})$.
@@
Let the setting be as in Theorem 4.5, with $\mathcal{M}$ given by equation (9). Let $\mathcal{H}_{V}$ be the Hecke operator attached to $\widetilde{\mathbf{X}}$, viewed as a Hecke polynomial via (11). Then there exists a collection of elements
@@
whenever $\mathfrak{m},\mathfrak{m}\ell\in\mathscr{R}$. Here, $\mathrm{Frob}_{\ell}^{-1}$ acts by its image in $\operatorname{Gal}(E[\mathfrak{m}]/E)$.
@@
    Projection to Galois representations: Kottwitz’s conjecture gives a concrete description of the cohomology group $\mathrm{H}^{d-1}(\mathrm{Sh}_{\mathbf{G}/\bar{E}},\mathbb{L}\otimes_{\mathbb{Z}_{p}}\mathbb{Q}_{p})$. In the case considered in Corollary 1.7, it gives a projection from this group to the Galois representation $\rho_{\pi}$. Moreover, in the cases considered below, the Hecke polynomial specialized to the Satake parameters of $\pi$ agrees with the characteristic polynomial of $\rho_{\pi}$.
@@ 4.4. Special cycles
Under the above assumptions, the discussions of [LS24, §2.4] carries over verbatim, and Proposition 2.7 of op. cit. gives a weaker version of the motivic theta series.
@@
Let $d=\dim\mathrm{Sh}_{\mathbf{H}}+1$. Let $\mathbb{L}$ be the $\mathbb{Z}_{p}$-local system on $\mathrm{Sh}_{\mathbf{G}}$ attached to $\xi$. The diagonal cycle construction defines a map
$\Theta:C_{c}^{\infty}(\mathbf{H}(\mathbb{A}^{p\infty})\backslash\mathbf{G}(\mathbb{A}^{p\infty}),\mathbb{Z}_{p})\to\mathrm{H}_{\mathrm{cont}}^{2d}(\mathrm{Sh}_{\mathbf{G}},\mathbb{L}(d))$
@@
  1. (1)
    The reason this is weaker than a motivic theta series is that for a general ring $R$, it is not true that
    $\mathbf{H}(R)\backslash\mathbf{G}(R)=\mathbf{X}(R).$
    This holds in the Gan–Gross–Prasad case below, but it fails for the unitary Friedberg–Jacquet case. This failure is related to the issue of stabilizing the relative trace formula.
    However, this is enough for the purpose of constructing Euler systems. In fact, our main local result Proposition 3.18 gives an abstract tame norm relation for each $\mathbf{G}(F_{\ell})$-orbit of $\mathbf{X}(F_{\ell})$.
  2. (2)
@@
We now explain the content of Corollary 4.6 in a few cases. In each case, we need to specify the data of an embedding of reductive groups $\iota:\mathbf{H}\hookrightarrow\mathbf{G}$ and a character $\nu:\mathbf{H}\to\mathrm{U}(1)$. The latter is used to specify the $\mathrm{U}(1)$-bundle over $\mathbf{X}$. In most cases, the local spherical variety appears in the table [Sak13, Appendix A], so they have been verified to satisfy the combinatorial conditions alluded to in §3.3.2. The exception is the twisted Friedberg–Jacquet case, where we derive analogoues of Sakellaridis’s results ourselves.
@@ 4.4.1. Gan–Gross–Prasad [LS24]
The intersection of $\mathbf{H}$ with a good Borel $\mathbf{B}$ is trivial, so the bundle is automatically combinatorially trivial. The Hecke operator attached to $\mathbf{X}$ corresponds to the standard tensor product representation of degree $n(n+1)$, so Corollary 4.6 recovers [LS24, Proposition 5.4], bypassing the explicit matrix calculations in §4 of op. cit.
@@ 4.4.2. Friedberg–Jacquet [GS23]
Let $\mathtt{V}_{2n}$ be a Hermitian space with signature (1,2n1)(1,2n-1) at one archimedean place and $(0,2n)$ at the other archimedean places. Let $\mathtt{W}$ be a totally definite subspace of dimension $n$, and let $\mathtt{W}^{\perp}$ be its dual. Consider the setting
@@
[…]
@@
The Hecke operator attached to $\mathbf{X}$ corresponds to the standard representation of degree $2n$, so Corollary 4.6 recovers the tame part of the split anticyclotomic Euler system for symplectic representations constructed by Graham and Shah [GS23, Proposition 9.27] without assumptions on the prime pp.
4.4.3. Twisted Friedberg–Jacquet
We now consider the global setting considered in [LXZ25], so we will change notations to align with theirs. To orient the reader, our Euler system will be a split anticyclotomic Euler system defined relative to the CM extension $F^{\prime}/F_{0}$ in the notations below.
We place ourselves in the setting considered in §1.2 from the introduction, which we briefly summarize. Recall that $E/F_{0}$ is a biquadratic extension with the following intermediate subfields
[…]
Moreover, $(B,\ast)$ is a division algebra over $F$ of dimension $(2n)^{2}$ with an embedding $E\hookrightarrow B$. They give rise to the algebraic groups
$\mathbf{H}=\operatorname{Res}_{E_{0}/F_{0}}\mathrm{U}(B^{E})=\{g\in(B^{E})^{\times}|g^{*}g=1\},\quad\mathbf{G}=\mathrm{U}(B)=\{g\in B^{\times}|g^{*}g=1\}.$
Both groups have Shimura varieties with reflex fields (contained in) $F$ and dimensions
$\dim_{F}\mathrm{Sh}_{\mathbf{H}}=n-1,\quad\dim_{F}\mathrm{Sh}_{\mathbf{G}}=2n-1.$
Let $\mathbf{T}=\mathrm{U}(1)_{F^{\prime}/F_{0}}$. The reduced form defines a combinatorially trivial character $\nu:\mathbf{H}\to\mathbf{T}$, and we have the augmented embedding
$\mathbf{H}\hookrightarrow\widetilde{\mathbf{G}}=\mathbf{G}\times\mathbf{T},\quad h\mapsto(h,\nu(h))$
After base changing to the algebraic closure, this and the previous untwisted case agree.
Let $\ell$ be a place of $F_{0}$ which splits in $F^{\prime}$. Away from finitely many bad places, there are two cases:
  • $\ell$ splits in $F$: the local calculation is explained in §3.4.2.
  • $\ell$ is inert in $F$: the local calculation is explained in §3.4.3
Therefore, we can apply Corollary 4.6 to obtain a motivic Euler system in this setting. We will now unwind the relevant definitions and describe the resulting Euler system.
Recall the following automorphic context from [LXZ25]. Let $\pi_{0}$ be a symplectic automorphic representation of $\mathrm{GL}_{2n}(\mathbb{A}_{F_{0}})$ whose base change to $F$ is cuspidal. Let $\pi$ be the automorphic representation of $\mathbf{G}(\mathbb{A}_{F_{0}})$ such that $\pi_{0}$ and $\pi$ have the same base change to $F$. There is a factorization
(12) $L(s,\mathrm{BC}_{E}(\pi_{0}))=L(s,\mathrm{BC}_{F^{\prime}}(\pi_{0}))L(s,\mathrm{BC}_{F^{\prime}}(\pi_{0}\otimes\eta_{F/F_{0}}))$
Stated imprecisely, [LXZ25, Conjecture 1.1] relates the non-vanishing of the $\mathbf{H}$-period for $\pi$ to the non-vanishing of one of the two $L$-functions on the right hand side.
Let $\rho_{\pi}$ be the Galois representation attached to $\pi_{0}$. The Kottwitz conjecture (cf. [BR94, Conjecture 5.2]) predicts that $\rho_{\pi}|_{\operatorname{Gal}_{F}}$ appears in the $\pi$-isotypic component of the cohomology of $\mathrm{Sh}_{\mathbf{G}}$. Let $\chi$ be an anticyclotomic Hecke character of $\mathbb{A}_{F^{\prime}}^{\times}$. Following the Abel–Jacobi map procedure described in §4.3, our motivic Euler system gives rise to a class in
$\mathrm{H}^{1}(E,\operatorname{Res}^{E}_{F}\rho_{\pi}\otimes\operatorname{Res}^{E}_{F^{\prime}}\chi)=\mathrm{H}^{1}(F^{\prime},(\operatorname{Ind}_{F^{\prime}}^{E}\operatorname{Res}^{E}_{F}\rho_{\pi})\otimes\chi),$
where the equality is a consequence of Shapiro’s lemma and the projection formula. We have a decomposition
$\operatorname{Ind}_{F^{\prime}}^{E}\operatorname{Res}^{E}_{F}\rho_{\pi}=\rho_{\pi}|_{\operatorname{Gal}_{F^{\prime}}}\oplus(\rho_{\pi}\otimes\eta_{F/F_{0}})|_{\operatorname{Gal}_{F^{\prime}}}.$
Therefore, we have constructed Selmer classes for a decomposable representation, whose pieces exactly correspond to the factorization (12) on the automorphic side.
The characteristic polynomial in this case corresponds to the local $L$-factor for $L(s,\mathrm{BC}_{E}(\pi_{0}))$. This is exactly the Hecke operator used in the non-split case (Proposition 3.16). In the split case, it is the square of the Hecke operator used in Proposition 3.14, since $\eta_{F/F_{0}}$ is trivial. This does not change the main result, as explained in Remark 3.19. Therefore, Corollary 4.6 gives a motivic Euler system, and the discussions thereafter gives Theorem 1.3 from the introduction. Note that in this case, Kottwitz’s conjecture is proven in [KSZ21].
Corollary 4.10.
Let $\pi$ be as above. Suppose Kottwitz’s conjecture holds for the $\pi$-isotypic part of the cohomology of $\mathrm{Sh}_{\mathbf{G}}$, then there exists a split anticyclotomic Euler system for the decomposable representation
$(\rho_{\pi}\oplus\rho_{\pi}\eta_{F/F_{0}})|_{\operatorname{Gal}_{F^{\prime}}}$
whose base class is the $p$-adic étale realization of the special cycle corresponding to the embedding $\mathrm{Sh}_{\mathbf{H}}\hookrightarrow\mathrm{Sh}_{\mathbf{G}}$.
@@ 4.5. Eisenstein classes
If $\mathbf{H}=\mathrm{GL}_{2}$ and $V$ is its standard two-dimensional representation, then Example 4.3 gives a $V$-theta series as above, with $i=j=1$. This gives rise to the tame parts of the following Euler systems, with the caveat that we need the function-level statements of [SW22] in mixed characteristics. Note that all of the spherical varieties used are strongly tempered.